OpenStudy (anonymous):

Find the function using f(x)= x+1/x g(x)= x+1/x+2 f o f

sam (.sam.):

\[ f(x)=\frac{ x+1}{x}\]?

OpenStudy (anonymous):

its actually x^4+3x^2+1/x(x^2+1) i'm juts not sure how they got that

sam (.sam.):

Yes, you function is \[f(x)=x+\frac{1}{x}\] \[f o f=f(f(x))\\=f(x+\frac{1}{x})\\=(x+\frac{1}{x})+\frac{1}{(x+\frac{1}{x})}\]

sam (.sam.):

\[=x+\frac{1}{x}+\frac{x}{x^2+1}\] \[=\frac{x^2+1}{x}+\frac{x}{x^2+1}\] Common denominators and you should get what you wanted

OpenStudy (anonymous):

ok thank you

sam (.sam.):

Do you understand? You can ask if you don't :)

OpenStudy (anonymous):